A meeting of consuls was attended by 4 Americans and 2 Germans. If three consuls were selected at random one after the other, illustrate the probability distribution of a random variable G (number of Germans) and draw the histogram.

1
Expert's answer
2022-03-03T12:32:42-0500

Let's denote G - Germans, A - Americans.

Sample space S is all possible outcomes.


"S = \{AAA, AAG, AGA, GAA, AGG, GAG, GGA\}"

Let "g" be the random variables representing the number of Germans that occur. The possible values of "g" are "0,1,2."



"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ point & g \\ \hline
 AAA & 0 \\
 \hdashline
 AAG & 1 \\
 \hdashline
 AGA & 1 \\
 \hdashline
 GAA & 1 \\
 \hdashline
 AGG & 2 \\
 \hdashline
 GAG & 2 \\
 \hdashline
 GGA & 2 \\
 \hdashline
\end{array}"


"P(G=0)=\dfrac{4}{6}(\dfrac{3}{5})(\dfrac{2}{4})=\dfrac{1}{5}"

"P(G=1)=\dfrac{4}{6}(\dfrac{3}{5})(\dfrac{2}{4})+\dfrac{4}{6}(\dfrac{2}{5})(\dfrac{3}{4})+\dfrac{2}{6}(\dfrac{4}{5})(\dfrac{3}{4})=\dfrac{3}{5}"

"P(G=2)=\dfrac{4}{6}(\dfrac{2}{5})(\dfrac{1}{4})+\dfrac{2}{6}(\dfrac{4}{5})(\dfrac{1}{4})+\dfrac{2}{6}(\dfrac{1}{5})(\dfrac{4}{4})=\dfrac{1}{5}"

The probability distribution of a random variable "G"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 g & 0 & 1 & 2 \\ \hline
 p(g) & 1/5 & 3/5 & 1/5
\end{array}"